OpenStudy (anonymous):

Research indicates that in a controlled environment, the number of diseased mice will increase linearly each day after one of the mice in the cage is infected with a particular type of disease causing germ. There were 8 diseased mice 4days after the first exposure and 14 diseased mice after 6 days. (a) Write an equation that will give the number of diseased mice after any given number of days.

OpenStudy (anonymous):

this is asking for the equation of a line through the points (4,8) and (6,14)

OpenStudy (anonymous):

find the slope \[m=\frac{14-8}{6-4}=\frac{6}{2}=3\]

OpenStudy (anonymous):

then use the point slope formula \[y-8=3(x-4)\] \[y-8=3x-12\] \[y=3x-4\]

OpenStudy (anonymous):

Thankyou very much satellite73.